OpenStudy (anonymous):

HELPPPPP! For this right triangle shown, what is the sine of angle C? A) 15 7 B) 15 8 C) 17 7 Eliminate D) 17 8

OpenStudy (anonymous):

tha answers say A) the square root of 15 over 8 B) the square root of 15 over 8 C) the square root of 17 over 7 D) the square root of 17 over 8

OpenStudy (mathstudent55):

Use the Pythagorean theorem to find the length of the side AB. Then remember that \(\sin \theta = \frac{opp}{hyp} \).

OpenStudy (mathstudent55):

Also, check your choices bec the first two choices above are the same.

OpenStudy (anonymous):

i mean A should be the square root of 15 over 7

OpenStudy (anonymous):

did you get c ?

OpenStudy (mathstudent55):

No. First, what did you get for the length of AB?

OpenStudy (anonymous):

i dont understand it

OpenStudy (anonymous):

my brother is saying it should be c